# Towards Coordinate- and Dimension-Agnostic Machine Learning for Partial Differential Equations

### v1

@@ B.1 Spatial-Differentiation
[…]
@@
[…]
@@ B.2 Temporal-Integration
[…]
@@
Appendix C Equations in Coordinate-Explicit Forms
Barkley System: In coordinate-explicit form, the Barkley system (Equation 8) is expressed as
Modified Patlak-Keller-Segel System: In coordinate-explict form, the Modified Patlak-Keller-Segel System 9) is expressed as
Appendix D Training the Neural Networks
@@
at each location $x$ in space and point $t$ in time; see Fig. 1A. For the FitzHugh-Nagumo system (Section 3.1), we set $n_{\text{neu}}=64$, with the fields identified as $(\Psi_{1},\Psi_{2})=(V,W)$. In the Barkley system (Section 3.2), we likewise use $n_{\text{neu}}=64$, where $(\Psi_{1},\Psi_{2})=(U,V)$. With the Patlak-Keller-Segel system (Section 3.3), we consider $n_{\text{neu}}=32$, in which $(\Psi_{1},\Psi_{2})=(A,B)$.
The training data for each case is first simulated and then subsampled for training as described below: Section D.1 for the FitzHugh-Nagumo system, Section D.2 for the Barkley system, and Section D.3 for the modified Patlak-Keller-Segel system. For all simulations in this work, we have checked that the numerical computations are robust to space-time discretizations.
@@
D.1 FitzHugh-Nagumo System Training
@@
We generate numerical simulation data of the FitzHugh-Nagumo system on a one-dimensional domain specified in Eq. (30), using $L=10^{2}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-1},1)$, and for every numerical simulation of Eq. (7) we generate the time progression consecutively at each timestep $\Delta t$, using forward Euler method Eq. (27), for $t\in[0,T]$ where $T=10^{3}$. All training data is obtained from a single simulation, in which the initial field configuration — consisting of combinations of a few localized configurations — is given in Appendix E.1. Then, we randomly-selected a subset of the simulation data (i.e. $10\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer [28], set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
D.2 Barkley System Training
@@
We create numerical simulation data on the one-dimensional domain specified in Eq. (30) for Barkley system, using $L=5\times 10^{1}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-2},5\times 10^{-1})$, and for every numerical simulation of Eq. (8) we generate the time progression consecutively at each timestep $\Delta t$, using the forward Euler method Eq. (27), for $t\in[0,T]$ where $T=2\times 10^{2}$. All training data is obtained from a single simulation, in which the initial field configuration — consisting of combinations of a few localized configurations — is given in Appendix E.2. Then, we randomly-selected a subset of the simulation data (i.e. $5\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer, set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
D.3 Patlak-Keller-Segel System Training
@@
We produce numerical simulation data for the Palak-Keller-Segel system on a one-dimensional domain specified in Eq. (30), using $L=10^{2}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-1},1$, and for every numerical simulation of Eq. (9) we generate the time progression consecutively at each timestep $\Delta t$, using forward Euler method Eq. (27), for $t\in[0,T]$ where $T=5\times 10^{2}$. The training data is derived from multiple simulations showcasing traveling wave collisions and diffusion. The initial field configurations used to create these simulations are uniform for $A$ and consist of combinations of a few localized configurations for $B$, as provided in Appendix E.3. Our choice of domain dimensions and initial field configurations are motivated by their relevance to real experimental settings using microfluidic devices. Then, we randomly-selected a subset of the simulation data (i.e. $2\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer, set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
@@
Appendix E Initial Field Configurations for Training
Here we list the initial field configurations used to generate the training data for the FitzHugh-Nagumo, the Barkley, and the Patlak-Keller-Segel models, as explained in Appendix D.
E.1 FitzHugh-Nagumo system
@@ E.2 Barkley system
E.3 Patlak-Keller-Segel system
Here we consider an active matter system of many chemotactic bacteria that navigate in response to chemo-attractants, which they simultaneously deplete through their own consumption [49]. Two important phenomena emerge from bacterial collective chemotaxis: the formation of traveling bands of bacteria and the collisions of those bands, which then diffuse [50]. To generate a series of in silico experiments illustrating both phenomena, we begin with simple initial conditions: the Asp is uniform, and bacteria are inoculated at two ports, each modeled by a Gaussian distribution. We consider four different initial profiles for Asp concentration and four times three distinct initial profiles for the bacterial field, i.e.
@@ E.3 Patlak-Keller-Segel system
Appendix F Cartesian and Non-Cartesian Coordinates
Here we show the relation between different choices of coordinate systems used in Section 4.1 and Section 4.2, from the mapping between locations to the calculation of relevant differential forms.
F.1 Inverse Function and Illustration of $y$ vs $x$
@@
Figure 5: The Cartesian and non-Cartesian coordinate systems in one- and two-dimensional space. Here we use Eq. (11) and Eq. (12) with $\lambda=L=10^{2}$. (A) One-dimensional space, conversion between $x$ and $y$. (B) Two-dimensional space, conversion between $(x_{1},x_{2})$ and $(y_{1},y_{2})$. The color transitions linearly with $x_{1}$, following the Python Matplotlib brg colormap [23], shifting from blue to red to green.
F.2 Calculation of Differential-Forms in One-Dimensional Euclidean Space
Denote ${}^{\prime}=d/dy$ when taking the derivative of the function $f(y)$. The differential-form inputs (besides $\Psi_{1}$ and $\Psi_{2}$) used in the neural network as illustrated in Fig. 1B for the one-dimensional non-Cartesian case (Section 4.1) can be calculated as follows:
@@ F.2 Calculation of Differential-Forms in One-Dimensional Euclidean Space
F.3 Calculation of Differential-Forms in Two-Dimensional Euclidean Space
The differential-form inputs (besides $\Psi_{1}$ and $\Psi_{2}$) used in the neural network as illustrated in Fig. 1B for the two-dimensional non-Cartesian case (Section 4.2) can be calculated as follows:
@@ F.3 Calculation of Differential-Forms in Two-Dimensional Euclidean Space
Appendix G Comparison between Different Coordinate Systems
We illustrate the comparison between coordinate systems — using 1D Cartesian and non-Cartesian coordinates as an example — which naturally extends to higher dimensions and curved spaces. For a given dynamical system, we initialize a condition (distinct from the training set) in $x$-space with uniform $\Delta x$ discretization. We then compute its evolution via both the true PDE and the neural network model from Section D, showing strong agreement. The corresponding initial condition in $y$-space (with uniform $\Delta y$) is evolved using the neural network, then mapped back to $x$-space via Eq. (11) for direct comparison. These steps are summarized in Fig. 6. For a coordinate-free model, all these generated datasets should be consistent, which we confirm through our results.

### v2

@@ B.1 Spatial-Differentiation
$\partial_{x}M(x,t)\overset{\text{num}}{\approx}\frac{M(x+\Delta x)-M(x-\Delta x)}{2\Delta x}\ .$
@@
$\partial_{x}^{2}M(x,t)\overset{\text{num}}{\approx}\frac{M(x+\Delta x)-2M(x)+M(x-dx)}{\Delta x^{2}}\ .$
Note that $\partial_{x}^{2}M(x,t)\overset{\text{num}}{\approx}\partial_{x}\left[\partial_{x}M(x,t)\right]$ up to a small error of order $\mathcal{O}(\Delta x^{2})$. When dealing with spatial derivatives inside other spatial derivatives, it can be beneficial to split them out rather than applying the derivatives consecutively before using the numerical approximations, e.g.
$\partial_{x}\left[N(x,t)\partial_{x}M(x,t)\right]\rightarrow\partial_{x}N(x,t)\partial_{x}M(x,t)+N(x,t)\partial_{x}^{2}M(x,t)\overset{\text{num}}{\approx}.$..\ . (17)
@@ B.2 Temporal-Integration
$\partial_{t}M(x,t)\overset{\text{num}}{\approx}\frac{M(x,t+\Delta t)-M(x,t)}{\Delta t}\ ,$ (20)
@@
Appendix C Equations in Coordinate-Free Form
Here we give a coordinate-free expression for the PDEs from Section 3.1 using the formalism from Appendix A.
FitzHugh-Nagumo System:
Barkley System:
Modified Patlak-Keller-Segel System:
Appendix D Numerical Results for Time-Evolution PDEs
The learning of all PDE systems is conducted on data from a one-dimensional periodic domain $x\in[-L,+L]$. The data and neural network architectures are detailed in Appendix E. All simulations use evenly spaced grids for finite-difference computations in the specified coordinate system, and we verify that the numerical schemes are robust to space-time discretizations (see Appendix I). For notational convenience, we denote the fields obtained by numerically integrating the exact PDE as $\Psi_{i}$, and the predictions generated by the neural network models as $\hat{\Psi}_{i}$. The argument $\eta(\zeta)$ denotes that the computation was performed in $\zeta$-coordinates and has been push-forwarded into $\eta$-coordinates via a coordinate transformation. A uniform grid in $\zeta$-coordinates will in general correspond to a non-uniform grid in $\eta$-coordinates.
All quantitative comparisons between the “true” and learned dynamics are contained in Appendix K.
D.1 One-Dimensional Euclidean Space
$x=f(y)\equiv x(y)=(y/2)\left[1+(y/\lambda)^{2}\right]\ ,$ (25)
as illustrated in Fig.2A. We set $\lambda=L$. The inverse function $y=f^{-1}(x)\equiv y(x)$ is provided in Appendix G.1. The input pre-processing for the neural network (Fig. 1B, in $y$-coordinate) is detailed in Appendix G.2.
We test the trained model on a new initial condition $t=0$ with uniform $A$ and localized peaks in $B$ (see Appendix I.1). The true dynamics, $A(x,t)$ and $B(x,t)$, are obtained by integrating Eq.(6) (Fig.2B1); the learned dynamics, $\hat{A}(x,t)$ and $\hat{B}(x,t)$, are generated by the neural network (Fig.2B2) are a good approximation. To test our coordinate-free generalization claims, we interpolate the initial field configurations to $y$-coordinates (see Appendix H), integrate the neural network model to get $\hat{A}(y,t)$ and $\hat{B}(y,t)$ (Fig.2D1), and then map the results back to $x$-space to arrive at $\hat{A}(x(y),t)$ and $\hat{B}(x(y),t)$ (Fig.2D2). We can also transform Fig. 1B1 to $y$-coordinates, obtaining $A(y(x),t)$ and $B(y(x),t)$ as shown in Fig. 2C. We observe that Fig. 2C and Fig. 2D1 agree, as well as Fig. 2A1 and Fig. 2D2, demonstrating that our machine learning method generalizes across coordinate systems.
D.2 Two- and Three-Dimensional Euclidean Space
$x_{1}=f(y_{1})\equiv x_{1}(y_{1})\ ,\ x_{2}=f(y_{2})\equiv x_{2}(y_{2})\ ,$ (26)
where $f$ is as in Eq.(25) and we set $\lambda=L$. The domain is the square $(x_{1},x_{2})\in[-L,+L]\times[-L,+L]$ with periodic boundaries. An illustration of $(y_{1},y_{2})$ versus $(x_{1},x_{2})$ is in Appendix G.1; input pre-processing in $(y_{1},y_{2})$-coordinates for the neural network is detailed in Appendix G.3.
Spiral waves can emerge for certain instances of the FitzHugh-Nagumo system. We choose an initial field configuration at $t=0$ where the spiral wave pattern is clearly visible (an explanation of how we do it is provided in Appendix I.2). In $(x_{1},x_{2})$-coordinates, the true dynamics are computed by integrating Eq.(4) (Fig. 3A1-4 top row, $V(x_{1},x_{2},t)$ and $W(x_{1},x_{2},t)$). In $(y_{1},y_{2})$-coordinates, the learned model is integrated from the extrapolated initial condition (see Appendix H for details), and the results are mapped back to $(x_{1},x_{2})$ (Fig. 3A1-4 bottom row, $\hat{V}(x_{1}(y_{1}),x_{2}(y_{2}),t)$ and $\hat{W}(x_{1}(y_{1}),x_{2}(y_{2}),t)$). We observe that the predictions agree, suggesting that our machine learning method not only generalizes across coordinate systems, but also it extends across dimensions. The correct predictions are made in two dimensions, even though the model was trained in one dimension, indicating that the approach is dimension-free.
$\star d\Psi_{i}\big|_{\partial\mathcal{C}}=0\ ,$ (27)
We initialize the Barkley system with a field configuration that clearly displays scroll wave patterns (Appendix I.3). Fig. 3B1–4 shows that the neural network predictions, $\hat{U}(x_{1}(y_{1}),x_{2}(y_{2}),x_{3}(y_{3}),t)$ and $\hat{V}(x_{1}(y_{1}),x_{2}(y_{2}),x_{3}(y_{3}),t)$, closely match the true dynamics, $U(x_{1},x_{2},x_{3},t)$ and $V(x_{1},x_{2},x_{3},t)$, as governed by Eq. (5), demonstrating the adaptability of our dimension-free model to diverse boundary conditions.
D.3 Two-Dimensional Curved Surface Embedded in 3-Dimensional Euclidean Space
In addition to increasing the number of the spatial dimensions and modifying the boundary conditions, we now bend the space (intrinsically). For this case, we explore the dynamics of the FitzHugh–Nagumo fields on a two-dimensional domain (similar to Fig. 3A but not flat) embedded in a spherical surface — a manifold with constant positive intrinsic curvature. We consider two coordinate representations of this curved space — the geographic $(\theta,\phi)$ and the stereographic $(\rho,\varphi)$ systems — as illustrated in Fig. 6A-C (See Appendix J for more details). The transformation between these coordinate systems is given by:
$\rho=R\tan\left(\phi/2\right)\ ,\ \varphi=\phi\ ,$ (28)
The spatial domain will be constrained between the radial values $\rho_{\min}$ and $\rho_{\max}$, and the angular values $\varphi_{\min}$ and $\varphi_{\max}$, which can be determined from $\theta_{\min}$, $\theta_{\max}$, $\phi_{\min}$, $\phi_{\max}$ (Appendix I.4), We impose heterogeneous boundary conditions: the Neumann (no-flux) conditions on the latitude boundaries and the periodic conditions on the longitude boundaries.
We initialize the system with a clearly defined spiral pattern for the fields $V$ and $W$ (defined in $(\theta,\phi)$-coordinates, see Appendix I.4). The true dynamics — $V(x(\theta,\phi),y(\theta,\phi),z(\theta,\phi),t)$ and $W(x(\theta,\phi),y(\theta,\phi),z(\theta,\phi),t)$ — are obtained by integrating Eq.(4) in $(\theta,\phi)$-coordinates and then mapped to the embedding space $(x,y,z)$, as illustrated in the top row of Fig.6D1–4. For the stereographic $(\rho,\varphi)$-coordinates, we extrapolate the initial condition (Appendix H), integrate the neural network learned dynamics, and transform the results to $(x,y,z)$ to obtain $\hat{V}(x(\rho,\varphi),y(\rho,\varphi),z(\rho,\varphi),t)$ and $\hat{W}(x(\rho,\varphi),y(\rho,\varphi),z(\rho,\varphi),t)$, shown in the bottom row of Fig.6D1–4. We observe a remarkable agreement (quantified in Appendix K) between both predictions, illustrating that our machine learning method not only generalizes across coordinate systems, extends across dimensions, and adapts to different boundary conditions, but also remains robust under curvatures.
D.4 Two-Dimensional Flat Surface with Non-Trivial Topology
Finally, we change the topology of the environment. In essence, this is straightforward, since only the boundary specification needs to be modified. For example, let us study the FitzHugh–Nagumo system on a flat annulus in the polar coordinates $(\rho,\theta)$, which can be related to the Cartesian coordinates $(x,y)$ via:
$x=\rho\cos\theta\ ,\ y=\rho\sin\theta\ ,$ (30)
where the region of interests is confined within $\rho\in[\rho_{\min},\rho_{\max}]$, where we impose fixed-value Dirichlet boundary conditions.
Figure 7: The time evolution of the FitzHugh–Nagumo system on a two-dimension flat surface with an annulus topology. We present the time evolution of the true dynamics obtained by integrating Eq. (4) and compare it with the dynamics generated using the learned model at four different time frames: (A) $t=0$, (B) $t=15$, (C) $t=30$, (D) $t=45$. Each time frame consists of four two-dimensional plots: the top row shows the true dynamics, while the bottom row displays the neural network predictions; the left column corresponds to $V$-field and the right column to $W$-field of the FitzHugh-Nagumo system.
Similar to what has been done in Section D.2, we choose an initial field configuration at $t=0$ where the spiral wave pattern is clearly visible (an explanation of how we do it is provided in Appendix I.5). The true dynamics are computed by integrating Eq.(4) (Fig. 7A-D top row, $V(\rho,\theta,t)$ and $W(\rho,\theta,t)$). The learned model is integrated from the very same initial condition (Fig. 7A-D bottom row, $\hat{V}(\rho,\theta,t)$ and $\hat{W}(\rho,\theta,t)$). Both of these are mapped back to the $(x,y)$ coordinates, i.e. $V(x(\rho,\theta),y(\rho,\theta),t)$, $W(x(\rho,\theta),y(\rho,\theta),t)$, $\hat{V}(x(\rho,\theta),y(\rho,\theta),t)$, and $\hat{W}(x(\rho,\theta),y(\rho,\theta),t)$, for illustration purpose. We observe that the predictions agree, suggesting that our machine learning method not only generalizes across coordinate and intrinsically curved systems, but also applicable to non-trivial topologies. In other words, our “spatially liberated” approach is topology-free.
Appendix E Training the Neural Networks
@@
at each location $x$ in space and point $t$ in time; see Fig. 1A. For the FitzHugh-Nagumo system, we set $n_{\text{neu}}=64$, with the fields identified as $(\Psi_{1},\Psi_{2})=(V,W)$. In the Barkley system, we likewise use $n_{\text{neu}}=64$, where $(\Psi_{1},\Psi_{2})=(U,V)$. With the Patlak-Keller-Segel system, we consider $n_{\text{neu}}=32$, in which $(\Psi_{1},\Psi_{2})=(A,B)$.
The training data for each case is first simulated and then subsampled for training as described below: Section E.1 for the FitzHugh-Nagumo system, Section E.2 for the Barkley system, and Section E.3 for the modified Patlak-Keller-Segel system. For all simulations in this work, we have checked that the numerical computations are robust to space-time discretizations.
@@
E.1 FitzHugh-Nagumo System Training
@@
We generate numerical simulation data of the FitzHugh-Nagumo system on a one-dimensional domain specified in Eq. (31), using $L=10^{2}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-1},1)$, and for every numerical simulation of Eq. (22) we generate the time progression consecutively at each timestep $\Delta t$, using forward Euler method Eq. (21), for $t\in[0,T]$ where $T=10^{3}$. All training data is obtained from a single simulation, in which the initial field configuration — consisting of combinations of a few localized configurations — is given in Appendix F.1. Then, we randomly-selected a subset of the simulation data (i.e. $10\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer Kingma (2014), set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
E.2 Barkley System Training
@@
We create numerical simulation data on the one-dimensional domain specified in Eq. (31) for Barkley system, using $L=5\times 10^{1}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-2},5\times 10^{-1})$, and for every numerical simulation of Eq. (23) we generate the time progression consecutively at each timestep $\Delta t$, using the forward Euler method Eq. (21), for $t\in[0,T]$ where $T=2\times 10^{2}$. All training data is obtained from a single simulation, in which the initial field configuration — consisting of combinations of a few localized configurations — is given in Appendix F.2. Then, we randomly-selected a subset of the simulation data (i.e. $5\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer, set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
E.3 Patlak-Keller-Segel System Training
@@
We produce numerical simulation data for the Palak-Keller-Segel system on a one-dimensional domain specified in Eq. (31), using $L=10^{2}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-1},1$, and for every numerical simulation of Eq. (24) we generate the time progression consecutively at each timestep $\Delta t$, using forward Euler method Eq. (21), for $t\in[0,T]$ where $T=5\times 10^{2}$. The training data is derived from multiple simulations showcasing traveling wave collisions and diffusion. The initial field configurations used to create these simulations are uniform for $A$ and consist of combinations of a few localized configurations for $B$, as provided in Appendix F.3. Our choice of domain dimensions and initial field configurations are motivated by their relevance to real experimental settings using microfluidic devices. Then, we randomly-selected a subset of the simulation data (i.e. $2\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer, set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
@@
Appendix F Initial Field Configurations for Training
Here we list the initial field configurations used to generate the training data for the FitzHugh-Nagumo, the Barkley, and the Patlak-Keller-Segel models, as explained in Appendix E.
F.1 FitzHugh-Nagumo system
@@ F.2 Barkley system
F.3 Patlak-Keller-Segel system
Here we consider an active matter system of many chemotactic bacteria that navigate in response to chemo-attractants, which they simultaneously deplete through their own consumption Phan et al. (2024). Two important phenomena emerge from bacterial collective chemotaxis: the formation of traveling bands of bacteria and the collisions of those bands, which then diffuse Phan et al. (2020). To generate a series of in silico experiments illustrating both phenomena, we begin with simple initial conditions: the Asp is uniform, and bacteria are inoculated at two ports, each modeled by a Gaussian distribution. We consider four different initial profiles for Asp concentration and four times three distinct initial profiles for the bacterial field, i.e.
@@ F.3 Patlak-Keller-Segel system
Appendix G Cartesian and Non-Cartesian Coordinates
Here we show the relation between different choices of coordinate systems used in Section D.1 and Section D.2, from the mapping between locations to the calculation of relevant differential forms.
G.1 Inverse Function and Illustration of $y$ vs $x$
@@
Figure 8: The Cartesian and non-Cartesian coordinate systems in one- and two-dimensional space. Here we use Eq. (25) and Eq. (26) with $\lambda=L=10^{2}$. (A) One-dimensional space, conversion between $x$ and $y$. (B) Two-dimensional space, conversion between $(x_{1},x_{2})$ and $(y_{1},y_{2})$. The color transitions linearly with $x_{1}$, following the Python Matplotlib brg colormap Hunter (2007), shifting from blue to red to green.
G.2 Calculation of Differential-Forms in One-Dimensional Euclidean Space
Denote $^{\prime}=d/dy$ when taking the derivative of the function $f(y)$. The differential-form inputs (besides $\Psi_{1}$ and $\Psi_{2}$) used in the neural network as illustrated in Fig. 1B for the one-dimensional non-Cartesian case (Section D.1) can be calculated as follows:
@@ G.2 Calculation of Differential-Forms in One-Dimensional Euclidean Space
G.3 Calculation of Differential-Forms in Two-Dimensional Euclidean Space
The differential-form inputs (besides $\Psi_{1}$ and $\Psi_{2}$) used in the neural network as illustrated in Fig. 1B for the two-dimensional non-Cartesian case (Section D.2) can be calculated as follows:
@@ G.3 Calculation of Differential-Forms in Two-Dimensional Euclidean Space
Appendix H Comparison between Different Coordinate Systems
We illustrate the comparison between coordinate systems — using 1D Cartesian and non-Cartesian coordinates as an example — which naturally extends to higher dimensions and curved spaces. For a given dynamical system, we initialize a condition (distinct from the training set) in $x$-space with uniform $\Delta x$ discretization. We then compute its evolution via both the true PDE and the neural network model from Section E, showing strong agreement. The corresponding initial condition in $y$-space (with uniform $\Delta y$) is evolved using the neural network, then mapped back to $x$-space via Eq. (25) for direct comparison. These steps are summarized in Fig. 9. For a coordinate-free model, all these generated datasets should be consistent, which we confirm through our results.